na2006x:

1,200 employees of a company were surveyed to find out whether they were satisfied with the company’s insurance policy. The survey showed that 80% of the respondents were not satisfied with the policy. The standard error of the proportion is %. The number of people who are satisfied with the policy is between % and %.

mintyy:

what are you asking for?

Luigi0210:

You're missing some information. The standard error is defined as \(\Large SE = \frac{\sigma}{\sqrt{n}} \) But in order to find \(\sigma\) we need the original data (x) and the mean. \(\large \sigma = \sqrt{\frac{\sum (x_i-x)^2}{n-1}} \)

CaspianWilde:

the standard error of the proportion is like 1.16%.

mintyy:

@caspianwilde wrote:
the standard error of the proportion is like 1.16%.
how did u get that

CaspianWilde:

@mintyy wrote:
@caspianwilde wrote:
the standard error of the proportion is like 1.16%.
how did u get that
I did math!!! 😁